### v1

# Quantum reservoir computing for predicting and characterizing chaotic maps

Abstract
Quantum reservoir computing has emerged as a promising paradigm for harnessing quantum systems to process temporal data efficiently by bypassing the costly training of gradient-based learning methods. Here, we demonstrate the capability of this approach to predict and characterize chaotic dynamics in discrete nonlinear maps, exemplified through the logistic and Hénon maps. While achieving excellent predictive accuracy, we also demonstrate the optimization of training hyperparameters of the quantum reservoir based on the properties of the underlying map, thus paving the way for efficient forecasting with other discrete and continuous time-series data. Furthermore, the framework exhibits robustness against decoherence when trained in situ and shows insensitivity to reservoir Hamiltonian variations. These results highlight quantum reservoir computing as a scalable and noise-resilient tool for modeling complex dynamical systems, with immediate applicability in near-term quantum hardware.
Introduction.– Quantum algorithms with proven quantum advantage, such as prime number factorization [1, 2] and search [3, 4], are not easily implementable in near-term Noisy Intermediate Scale Quantum (NISQ) computers [5]. Thus, the quest for developing experiment-friendly heuristic quantum algorithms and identifying new applications is especially important in the current NISQ era. Variational quantum algorithms [6, 7, 8, 9, 10, 11] are one class of such algorithms which have attracted a lot of attention recently as a natural quantum extension of neural network-based learning models. However, they suffer from the same flaws as their classical counterparts - barren plateaus inhibiting training [12, 13, 14], as well as limitations in processing temporal correlations  [15]. Recently, quantum systems have been shown to achieve efficiencies beyond classical models for analyzing time series data [16]. In addition, inspired by advancements in classical echo state networks [17, 18], quantum reservoir computing (QRC) has emerged as an alternative paradigm of heuristic quantum algorithms, with the benefit of easier training and efficient capture of temporal correlations [19, 20, 21]. QRC has already found applications for time series analysis [19, 22, 23, 24], entanglement detection [25, 21], quantum tomography [26, 27], quantum estimation [28, 29] and quantum state preparation [30, 31] among others. QRC has also been proposed for various quantum platforms, such as photonic [25, 30, 32], coherently coupled quantum oscillators [33], neutral atomic Rydberg arrays [34], nuclear spin-based reservoirs [35], and superconducting qubit platforms [36]. While the prediction capacity of QRC has been demonstrated for time series [19, 37, 38, 39, 40, 41, 24], its full potential for the prediction and characterization of signals with inherent nonlinearity, in particular generated by chaotic behavior, is less explored.
Classical discrete non-linear systems, which describe various situations, exhibit rich dynamical behaviors [42]. These include sinks or stable fixed points, periodic orbits, limit cycles, and, strikingly, chaotic behavior where slight changes in initial conditions lead to divergent outcomes [43]. This is formalized through the notion of Lyapunov exponents, which characterizes the rate of separation of infinitesimally close trajectories. Formally, if two series $\{x_{t}\}$ and $\{x^{\prime}_{t}\}$ emanating from the same dynamical system with slight differences in the values of the control parameters are considered with initial separation $\delta_{0}{=}x_{0}{-}x^{\prime}_{0}$, the largest Lyapunov Exponent (LLE) $\lambda^{*}$ is defined as $\lambda^{*}{=}\lim_{t\to\infty,|\delta(0)|\to 0}(1/t)\ln{(|\delta(t)/\delta(0)|)}$, where $\delta_{t}{=}x_{t}{-}x^{\prime}_{t}$. If $\lambda^{*}{<}0$, the perturbations in the initial conditions decay exponentially, indicating that the system is stable. If $\lambda^{*}{=}0$, the system exhibits neutral stability commonly associated with periodic or quasi-periodic dynamics. In contrast, if $\lambda^{*}{>}0$, the trajectories originating from infinitely close initial conditions diverge exponentially over time, a hallmark of chaotic behavior. This makes the prediction of future outcomes of time series generated from chaotic maps a truly challenging task. Thus, several questions arise: (i) Can quantum machine learning algorithms, such as QRC, succeed in forecasting? and (ii) Can QRC setups be used for characterizing the underlying features, such as nonlinearity and memory?
Figure 1: Schematic of QRC architecture. Quantum reservoir comprises two subsystems: (i) input qubits (blue) for encoding input variables through Pauli $Y$-rotation $R_{Y}$, and (ii) hidden qubits (pink) storing historical information to predict future outcomes. This process begins at time step tdt{-}d and continues until t1t{-}1, such that the total number of quantum layers is $d$. After the evolutions of $d$ sequential layers, quantum measurements along Pauli $X$-directions are implement on all the input and hidden qubits.
In this Letter, we answer these questions in the affirmative by demonstrating that discrete nonlinear dynamical maps can indeed be learned by quantum reservoir computing, leading to accurate forecasts even in the presence of fully developed chaotic dynamics. Specifically, we demonstrate the success of a quantum reservoir consisting of linearly connected XY chains in predicting the behavior of two canonical nonlinear dynamical maps, viz., the logistic map [42] and the Hénon map [44]. Analyzing the predictive success based on resource consumption strategies such as increasing the number of quantum layers and different encoding repetitions yields valuable insights towards efficient generalization of our approach to arbitrary nonlinear systems. Furthermore, we demonstrate the robustness of our approach through assessing the impact of quantum decoherence noise as well as random choice of reservoir Hamiltonians on the prediction performance.
Chaotic maps.– We consider two canonical non-linear dynamical maps, viz. the logistic and the Hénon maps, which have one and two control parameters, respectively, for generating time series. The goal is to use a quantum reservoir algorithm for predicting and characterizing such time series. The logistic map is the one-dimensional recurrence relation
$x_{t}=rx_{t-1}(1{-}x_{t-1}),$ (1)
where $r$ is the control parameter generally taken between $r{\in}[0,4]$, and $x_{t}{\in}[0,1]$. The map stems from Verhulst’s logistic growth law of the 19th century, subsequently finding application as a discrete model of population dynamics  [42], and is now widely popular as a pedagogical minimal model to demonstrate chaotic properties, as varying a single parameter $r$ reveals fixed points, period-doubling cascades, windows of periodicity, and fully chaotic regimes  [43].
The Hénon map is a two-dimensional discrete map originally introduced to study strange attractors in dissipative systems  [44]. It is given by the two-dimensional recurrence relation
$x_{t}=1{-}ax_{t-1}^{2}{+}y_{t-1}\quad;\quad y_{t}=bx_{t-1}\quad.$ (2)
The Hénon map provides a minimal, analytically accessible model of deterministic chaos in low dimensions. With a=1.4a{=}1.4 and b=0.3b{=}0.3, it produces the famous Hénon attractor. The map represents the stretch-and-fold mechanism of chaos in compact polynomial form. Each iteration stretches one direction, folds the trajectories back and contracts the area by $|b|$, creating a fractal invariant set with sensitive dependence on the initial conditions [45].
Quantum reservoir computing.– In the quantum reservoir computing (QRC) framework for classical time-series forecasting [19, 46, 20, 47, 48, 49], qubits are encoded with the time-series data at various time steps, and then allowed to sequentially interact with a quantum reservoir, whose Hamiltonian remains fixed throughout the training and prediction processes. This sequential interaction scrambles the encoded information inside the reservoir, which is then decoded through a final set of measurements enabling us to make a prediction for future time-steps. Unlike the paradigm of variational quantum algorithms, where one seeks to iteratively update all the parameters of a parametrized quantum circuit to optimize a given loss function, the QRC approach employs training only once, exclusively at the final readout stage. This mirrors classical reservoir computing, with the difference that the reservoir is implemented through a quantum system rather than a classical one.
QRC protocol for chaotic map prediction.– Let us now elaborate our implementation of QRC as depicted in Fig. 1 step-by-step. The reservoir collectively consists of two components - (i) qubits encoding information about the dynamical map (blue qubits in Fig. 1), which are replaced after each layer with fresh qubits encoding updated information; and (ii) so-called hidden qubits (pink qubits in Fig. 1) which are not directly manipulated until the very end, when at the output layer, all qubits in the reservoir are measured locally. The QRC procedure is as follows. (i) Reservoir engineering– Our reservoir is a linearly connected transverse XY chain with Hamiltonian given by
$H{=}J\sum_{j=1}^{N}(X_{j}X_{j+1}{+}Y_{j}Y_{j+1}){+}\sum_{j}h_{j}Z_{j},$ (3)
where $\{X_{j},Y_{j},Z_{j}\}$ are respectively the Pauli operators acting on the $j$-th qubit. The exchange couplings $J$ are fixed to unity in this study, and the magnetic field strengths $\{h_{i}\}$ are randomly sampled from within the ordered phase $h_{i}{\in}[0,1]$, since in the disordered phase ($h_{i}{>}1$), the spins are effectively isolated and thus not useful for percolation of information into hidden qubits. (ii) Encoding– Each input data $\vec{x}_{i}$ is a segment of a time series $i$ of the form $\vec{x}_{i}{=}(x^{(i)}_{t-d},x^{(i)}_{t-d+1},\cdots,x^{(i)}_{t-1})$, generated from a chaotic map, in which the last $d$ values are used to predict the next step $x^{(i)}_{t}$. For encoding the data, each data point of $\vec{x}_{i}$ is fed, oldest first, into the reservoir through $d$ sequential layers, where each layer consists of one elements of $\vec{x}_{i}$ encoding into qubits through a Pauli Y-rotation. To introduce nonlinearity [50, 51], $n_{\textrm{rep}}$ repetitions of the encoded qubit may be deployed at each layer. After each layer, the reservoir evolves under Hamiltonian $H$ for a time $\tau$, following which the original encoding qubits are discarded (i.e. traced out) and replaced by fresh qubits encoding the next elements of $\vec{x}_{i}$ and the state of hidden qubits $\rho_{\textrm{H}}$ is maintained to the next layer. This process continues until the information at time step t1t{-}1 is fed into the quantum reservoir. (iii) Readout– After the final round of evolution following the encoding of $x^{(i)}_{t-1}$, all input and hidden qubits (the numbers are denoted as $n_{\textrm{I}}$ and $n_{\textrm{H}}$, respectively) in the reservoir are locally measured along Pauli X-direction to yield a vector of expectation values $\vec{m}_{i}{=}[\langle X_{1}\rangle,\langle X_{2}\rangle,\cdots,\langle X_{n_{\textrm{I}}+n_{\textrm{H}}}\rangle]^{T}$. (iv) Training and Prediction– The core of QRC is its simple prediction method which is based on linear regression. In this approach, $x^{(i)}_{t}$ is estimated through a linear map like $x^{(i)}_{t}{=}\textbf{W}\vec{m}_{i}$ in which the weight matrix W has to be trained. This training is repeated through rolling windows over $s$ different training data sets $\{\vec{x}_{1},\vec{x}_{2},\cdots,\vec{x}_{s}\}$, each yielding a different measurement outcome $\vec{m}_{i}$. By making a matrix of such measurement outcomes $\textbf{M}{=}[\vec{m}_{1},\vec{m}_{2},\cdots,\vec{m}_{s}]$ and considering the true values of the time series $\textbf{Y}{=}[x^{(1)}_{t},x^{(2)}_{t},\cdots,x^{(s)}_{t}$] one can train the weight matrix W by minimizing the mean square error between the predicted value of $\textbf{W}\vec{m}_{i}$ and the corresponding true value $x^{(i)}_{t}$, namely $\text{MSE}{=}1/s\sum_{i}^{s}(x_{t}^{(i)}-\textbf{W}\vec{m}_{i})^{2}$. This results in the optimal weight matrix W as
$\textbf{W}^{*}=\textbf{Y}\textbf{M}^{\top}(\textbf{M}\textbf{M}^{\top}+\epsilon I)^{-1},$ (4)
where $I$ is the identity matrix and $\epsilon$ is a small value (we take it as $10^{-8}$ in this study) to ensure the robustness of the numerical calculation for matrix inversion. Hence, for predicting the unseen data point $x_{t}$ based on the $d$ past known values $(x_{t-d},x_{t-d+1},\cdots,x_{t-1})$, these values are similarly fed into the reservoir layer-by-layer to yield the final outcome vector $\vec{m}$. The prediction $\hat{x}_{t}$ for $x_{t}$ is then given by $\hat{x}_{t}{=}\textbf{W}^{*}\vec{m}$.
Predicting chaotic time series.– We are now in a position to illustrate its performance for forecasting of the logistic and Hénon maps mentioned before. We restrict ourselves to time series generated with $r{\in}[0,4]$ for the logistic map and $a{\in}[1,1.4]$ for the Hénon map. For the training set, for any given control parameter $r$ or $a$, we consider $100$ different time series with $20$ elements each, namely $\{x^{(i)}_{t}\}_{t=1}^{20}$ with index $i$ going from $1$ to $100$ denoting each time series. Each time series is generated by a different random initial element $x^{(i)}_{1}$ and at the end we normalize all the time series with the largest element among all the $100$ time series to ensure that all the elements remain between 0 and $1$. For the test data, we use the same approach for generating $10$ unseen time series with new initializations, each containing $200$ time steps.
Figure 2: The bifurcation diagram of the logistic map (a) and the Hénon map (b). The largest Lyapunov exponent (dark line) of the logistic map (c) and the henon map (d). The regions with a white background indicate LLE values less than 0, while the grey background highlights regions where the LLE is greater than 0. Additionally, the RMSE of predicting the bifurcation diagrams have shown as the blue lines in (c) and (d).
For the logistic map, we consider a quantum reservoir with d=2d{=}2 layers and $n_{\textrm{H}}{=}4$ hidden qubits. Since the logistic map only depends on one variable, the number of input qubits will be equal to the number of repetitions, namely $n_{\textrm{I}}{=}n_{\textrm{rep}}$. We keep $n_{\textrm{I}}{=}n_{\textrm{rep}}{=}2$ for simulating the logistic maps. For the Hénon map, we consider a quantum reservoir with d=1d{=}1 layer and $n_{\textrm{H}}{=}3$ hidden qubits. Since this Hénon map has two input variables, then $n_{\textrm{I}}{=}2n_{\textrm{rep}}$. Here, we take $n_{\textrm{I}}{=}4$ input qubits (i.e. $n_{\textrm{rep}}{=}2$ repetitions for each input data). We use the rolling window method for training. In Figs. 2(a) and (b), for any given control parameters $r$ and $a$, we depict the long time predictions from t=151t{=}151 to t=200t{=}200 for one of the test time series which shows very accurate prediction (blue points) for our QRC protocol. The bifurcation diagrams in Figs. 2(a) and (b) clearly single out the different dynamical regimes of the maps showing remarkable agreement with the corresponding true values (red points). Note that $x_{\infty}$ in these figures is represented by all the predictions for this test dataset for times between t=151t={151} to t=200t={200}. As the figure shows, for certain values of the control parameter, the long time behavior shows convergence to either a single value or oscillations between a few discrete values, indicating non-chaotic behavior. On the other hand, for some other values of the control parameters, the long time behavior tends to cover a continuous part of the space, which is an evidence of chaotic behavior. Interestingly, the predictions by the QRC can precisely identify the islands of periodicity between the chaotic regimes for both maps. For a more in-depth analysis, from the prediction in all test sets, we plot the corresponding Root Mean Square Error (RMSE) averaged over all test datasets in Figs. 2(c) and (d) for the logistic and Hénon maps respectively, where we also plot the corresponding LLEs for reference. Two results immediately appear – (i) the prediction error is lowered wherever the LLE is negative (sweeping across map parameter ranges – average of RMSE prediction errors across chaotic regimes vs non-chaotic regimes is $5.7\times 10^{-5}$ vs $8\times 10^{-4}$ for logistic map, and $6.3\times 10^{-5}$ vs $3.6\times 10^{-5}$ for Hénon map), and (ii) even within the extended regions where the LLE is positive, a very strong correlation (Spearman rank correlation coefficient $r_{s}{=}0.97$ for the logistic map, and $r_{s}{=}0.87$ for the Hénon map) between LLE and prediction error is observed. Together, these results show that the prediction error primarily inherent the degree of unpredictability of the dynamics.
Characterizing chaotic time series. – We have already demonstrated that our QRC framework is highly effective in performing single-step predictions. However, this still leaves a major question unaddressed - can one taylor the hyperparameters involved in the QRC implementation protocol to maximize the prediction accuracy for specific nonlinear dynamical maps? Conversely, assuming such an optimization strategy for one member of a specific family of nonlinear dynamical maps is obtained, can this strategy inform the QRC optimization strategy for other members of that family of maps as well?
To answer these questions, we confine ourselves to two hyperparameters of QRC, namely the number of repetitions $n_{\textrm{rep}}$ with which encoded qubits are introduced at each layer as well as the number of layers, i.e., circuit depth. Moreover, we consider the following family $\Phi$ of polynomial maps $\Phi:x_{t}{=}\sum_{j}P_{k_{j}}(x_{t-j})$, where $P_{k_{j}}$ is a polynomial of degree $k_{j}$. Both the logistic and the Hénon maps are members of this family. For example, for the logistic map the only nonzero contribution is coming from j=1j{=}1 for which $k_{1}{=}2$. In Fig. 3(a) and Fig. 3(b), we plot the prediction accuracy as a function of the number of layers as well as repetitions $n_{\textrm{rep}}$ for the logistic and Hénon maps respectively. The results strongly indicate that increasing the number of quantum layers or repetitions is not monotonic with prediction accuracy. For example, for the logistic map, best predictive accuracy for $x_{t}$ is achieved with two quantum layers and $n_{\textrm{rep}}{=}2$, corresponding to the input sequence $\{x_{t-2},x_{t-2},x_{t-1},x_{t-1}\}$. To explain the result, one recalls that the repetitions serve to induce nonlinearity in the otherwise linear quantum circuit. Thus $x_{t}$ is a nonlinear function of degree 2 of $x_{t-1}$ as per both the logistic and Hénon maps, therefore the expected optimal number of repetitions $n_{\textrm{rep}}{=}2$. From this, we hypothesize that the optimal repetition strategy $n_{\textrm{rep}}^{*}$ for the family of maps $\Phi$ is given by the highest polynomial degree i.e., $k_{\max}=\max_{j}(k_{j})$ as per the definition of $\Phi$ above. To test this hypothesis, we take the same logistic or Hénon maps, but now attempt to predict $x_{t+1}$ in terms of only $x_{t-1}$ and elements before. We note that $x_{t+1}$ is a degree-4 function of $x_{t-1}$ in both cases, and are thus also members of the family $\Phi$ in which the only non-zero contribution comes from j=1j{=1} with $k_{1}{=}4$. For example, the logistic map is now explicitly given as $x_{t+1}{=}{-}r^{3}x_{t-1}^{4}{+}2r^{3}x_{t-1}^{3}{-}r^{2}(1+r)x_{t-1}^{2}{+}r^{2}x_{t-1}$, Thus, we empirically expect $n_{\textrm{rep}}{=}4$ to be optimal for predicting two steps ahead. Indeed, Fig. 3(c) and Fig. 3(d) explicitly confirm this prediction numerically for both logistic and Hénon maps and thus support our hypothesis. As a corollary, this hypothesis also entails that quantum resource requirements for predicting chaotic systems over multiple time steps may sharply diverge. One can arrive at this conclusion in two steps. First, one notes that for the problem of predicting $x_{t{+}s}$, i.e., s+1s{+}1-steps in advance, the original map of a member of $\Phi$ with highest polynomial degree $k_{\max}$ transforms into another member of $\Phi$ with highest polynomial degree $k_{\max}^{s+1}$. Now, our hypothesis indicates that optimal input repetitions $n_{\textrm{rep}}^{*}$ must scale as $k_{\max}^{s+1}$, i.e., exponentially with $s$. This shows that predicting distant future outcomes of chaotic maps requires exponential resources, in terms of number of qubits.
Figure 3: The effect of number of layers and repetitions on QRC for the logistic and Hénon map for (a) and (b) predicting $x_{t}$ and (c) and (d) predicting $x_{t+1}$ for $r=3.75$ ($a=1.35$). Darker hues indicate lower RMSE (better predictive accuracy). Lowest RMSE is for 22 layers for logistic map (11 layer for Hénon map) and $n_{\textrm{rep}}{=}2$ for predicting $x_{t}$ and $n_{\textrm{rep}}{=}4$ for predicting $x_{t+1}$ in both cases.
Robustness of forecasting.– So far, our results have assumed perfect unitary evolution, as well as randomized but fixed instances of reservoir Hamiltonian initialization. However, in the NISQ era, the former is a simplistic assumption as the presence of noise leads to non-unitary evolution in real setups. Moreover, the appeal of reservoir computing lies in the fact that the algorithm should work well regardless of the exact internal coupling parameters of the reservoir. Thus, for actual implementation, we must account for both these factors to benchmark the realistic predictive accuracy of our protocol. This analysis has been performed in the end matter below. Firstly, we show that by training the system with the measured data from a noisy reservoir we can reach significant robustness against dephasing. Secondly, our statistical analysis demonstrates that our protocol is robust against randomness in reservoir couplings.
Conclusion.– In this work, we have proposed a QRC algorithm for predicting and characterizing nonlinear dynamics in discrete maps with chaotic regimes. We have demonstrated the effectiveness of our method on the logistic and Hénon maps. While we can achieve accurate prediction about the future steps of our time series, we have also been able to optimize the QRC encoding protocol based on the features of nonlinear maps studied herein, paving the way for further optimization of the QRC architecure for experimental time series data. In addition, the proposed QRC algorithm is resilient against decoherence and reservoir randomness, making it suitable for near-term quantum devices. We note here that in contrast to other recent proposals of applying QRC to chaotic dynamics forecasting [52, 53] studying continuous dynamical evolution, we study discrete chaotic maps. Moreover, we use simple phase encoding as opposed to computationally costly amplitude encoding [52]; and employ a simple reservoir only requiring nearest neighbor XX interaction which is readily available in superconducting quantum simulators [54, 55, 56, 57], ion-trap systems [58, 59, 60] and optical lattices [61, 62, 30] as opposed to requiring complex reservoir engineering with fully-connected [53] or non-sparse Hamiltonians [52].
Acknowledgments.– AB acknowledges support from the National Natural Science Foundation of China (grants No. 12274059, No. 12574528 and No. 1251101297). LM gratefully acknowledges the support of the “Hundred Talents” program of the University of Electronic Science and Technology of China, the “Outstanding Young Talents Program (Overseas)” of the National Natural Science Foundation of China, and the talent programs of the Sichuan province and Chengdu municipality.
I End Matter
In this end matter, we discuss the robustness of the QRC protocol.
I.1 Robustness against decoherence
So far, we assumed a perfect unitary evolution of the reservoir. However, in practice, unwanted interaction with the environment is unavoidable. Such interaction generally leads to decoherence, i.e., gradual loss of quantum properties in a physical system after some time, which is the biggest challenge to physical implementation of quantum algorithms [63]. Dephasing is the most common decoherence effect whose origin is the presence of random site-dependent magnetic fields in the environment. The effect of such fields is to change the unitary evolution into a Lindblad master equation described by
$\frac{\mathrm{d}\rho}{\mathrm{d}\tau}=-i~[H,\rho]+\gamma\sum_{k}\left(Z_{k}\rho Z_{k}^{\dagger}-\rho\right).$ (5)
where $\gamma$ represents the strength of the dephasing [63]. Here, we evaluate the performance of the QRC framework under phase damping noise using two different approaches. In the first approach, we train the QRC in an ideal (noise-free) environment but test it in the presence of noise. However, in the second approach, training and testing the QRC are both performed under the same noise. As shown in Figs. 4(a) and (b), the predictive accuracy of the first model rapidly declines with noise strength $\gamma$ for both the logistic and the Hénon maps, respectively. This decline comes as no surprise as our measurement readout is along on the Pauli-X basis, thus the information retrieved from the measurements also decays exponentially as the coherence decreases. However, surprisingly, the second approach performs exceptionally well and maintains robustness in accurate prediction even under reasonably large noise, leading to orders of magnitude better performance for both logistic and Hénon maps. These results strongly indicate in situ training is preferable in the QRC framework. Physically, this stems from the fact the predictor trained in situ succeeds in learning both the noise parameter $\gamma$ as well as the time-series data encoded, while the predictor trained on a noise-free environment naturally fails to account for the noise parameter $\gamma$ as a background parameter for the test data.
Figure 4: The RMSE of QRC with decoherent noise strength $\gamma$ for (a) the logistic map, and (b) the Hénon map. Blue (red) lines indicate training without (with) decoherence.
Figure 5: Histogram for RMSE statistics of 1000 randomly chosen Hamiltonian reservoirs for (a) logistic map, and (b) Hénon map. Red lines are Poisson distribution fits for $n_{\textrm{bin}}{=}40$ bins.
I.2 Robustness against reservoir randomness
In this section, we discuss the robustness of our QRC protocol for predicting chaotic maps with respect to the choice of the parameters in the quantum reservoir. In general, the lack of sensitivity to such parameterization is a highly desirable trait for reservoir computing, where the explicit goal is to only adjust the weight matrix W at the end. To check this, we randomly selected $1000$ Hamiltonians for both the logistic and the Hénon maps and binning their respective predictive errors, represented by RMSEs, in the histograms of Fig. 5. The results indicate that the majority of Hamiltonians yield low RMSE, with only a small fraction exhibiting big errors. Moreover, in both cases, the histograms are Poissonian in nature, which means only a very small fraction of all the Hamiltonians are ‘problematic’ with the vast majority of randomly chosen Hamiltonians yielding close to maximum predictive accuracy. This is consistent with the measure concentration phenomenon encountered in locally entangled many-body systems such as the present reservoir based on disordered transverse XY spin chains, where most dynamical states are in fact concentrated within a small subspace of the overall Hilbert space  [64, 65, 66]. While an explicit quantification is beyond the scope of the present work, it appears randomly sampled Hamiltonians do typically arise from this subspace and the exceptions outside this subspace are rare, thus a Poissonian treatment is apt.
Code Availability. The Code used in this paper can be found here in the Github Repositories.